OpenStudy (anonymous):

Pq=3x qr=6x+4 find pr q is between p and r

OpenStudy (epikrebel):

If pq and qr are equivalent, then the answer is -4/3.

OpenStudy (anonymous):

Keep going

OpenStudy (epikrebel):

I should have specified that x=-4/3 the answer is -8.

OpenStudy (anonymous):

Were trying to find Pr not X

OpenStudy (anonymous):

Wrong

OpenStudy (anonymous):

Can you help me with my question? I really need help

OpenStudy (anonymous):

you can't get negative eight . Just do the problem for me because the numbers don't make any sense to the answer

OpenStudy (anonymous):

my question is A group of 4 friends ae playing golf. The total cost of the round of golf is $108. Each person in the group has the same coupon. The total cost of the round with the coupons is $76. How much is the coupon worth? Math

OpenStudy (anonymous):

Get out of my problem please

OpenStudy (anonymous):

Sorry im new at this

OpenStudy (anonymous):

Keep Going

OpenStudy (epikrebel):

All you gave me was the equation. You didn't tell if they were equivalent or anything. All you told me was q is in between p and r.

OpenStudy (anonymous):

thats everything you need to know if q is the mid point then therefore the 2 equations are equivalent once we find x we substitute it in for prs value of 14x -6 which makes no sense to the books answer

OpenStudy (epikrebel):

Now you tell me its the mid point. You just told me that q is in between p and r. That doesn't tell me their relative distance.

OpenStudy (anonymous):

Did you get the answer yet

OpenStudy (anonymous):

Lol. This is funny. @Andreweinstienium Since your question has been hijacked, start a new one. (And post the link here so people go there.)

OpenStudy (anonymous):

Who hijacked my question?

OpenStudy (anonymous):

@maddymam ... Oh wait, nevermind. You guys are still talking about this question.

OpenStudy (anonymous):

\[\Large {pq=3x \\ qr=6x+4 \\ p < q< r \\ pr=? }\] That's all correct, right? (Especially the third line?)

OpenStudy (anonymous):

pr=14x-6

OpenStudy (anonymous):

did you get the answer

OpenStudy (epikrebel):

Wow pr=14x-6. You could have given us that at the beginning. PR=22

OpenStudy (anonymous):

Solvve it now?

OpenStudy (anonymous):

Can you show me the process?

OpenStudy (epikrebel):

pq+qr=pr so 3x+6x+4=14x-6.

OpenStudy (anonymous):

Oh makes sense